Buy a U.S. T-bond for $1,100 that makes annual coupon payments at a rate of 10% until the bond matures 10 years from now, at which time it will pay you $1,000


1
Expert's answer
2021-04-15T07:19:56-0400

trading price=11000

face value=1000

rate=10

time =10 years

annual interest="interest\times face value"

"=\frac{10}{100}\times1000"

"=100"

current yield"=\frac{annual interest}{trading price}"

"=\frac{100}{1000}=0.09091"

"=0.09091\times 100"

"=9.091" %

"yield to maturity=\frac{annual interest payment+(face value/years to maturity)}{face value +current price/2)}"

"=\frac{100-(100/2)}{2100/2}"

"=\frac{90}{1050}"

"=0.08571"

"=0.08571\times 100"

"=8.571" %